OpenStudy (anonymous):

find the dimensions of a rectangular garden with perimeter 62 yd and area 220 yd^2

OpenStudy (anonymous):

2a+2b=62 ab=220 Solve

OpenStudy (anonymous):

im still confused. a=220-x?

OpenStudy (anonymous):

a=220/b also 2a=-2b+62 a=-b+31

OpenStudy (anonymous):

where you get x from?

OpenStudy (anonymous):

i meant b

OpenStudy (anonymous):

thats a*b=220 so you have to divide a or b on both sides to seperate them

OpenStudy (anonymous):

well i see, but where am i getting the dimentions -b+31 isnt really an answer

OpenStudy (anonymous):

solve by substitution

OpenStudy (anonymous):

but theres only one equation?

OpenStudy (anonymous):

2a+2b=62 ab=220

OpenStudy (anonymous):

You are told the perimeter, area, and that its a rectangle. so...

OpenStudy (anonymous):

i must be missing something

OpenStudy (anonymous):

isolate either a or b in one of the equations i gave you, then plug what X= into the OTHER equation. youll get a=number. then you plug that in and you can find b=

OpenStudy (anonymous):

y=1?

OpenStudy (anonymous):

i mean b

OpenStudy (anonymous):

have you learned how to solve systems of equations by substitution or elimination?

OpenStudy (anonymous):

well its what im learning now, so im just confused

OpenStudy (anonymous):

which two equations am i substituting? you solved before so im wondering

OpenStudy (anonymous):

2a+2b=62 is one ab=220 is the other 2(220/b)+2b=62 b=???? Put that into the ab=220 one to find a

OpenStudy (anonymous):

whats 2(220/b)?

OpenStudy (anonymous):

\[{2\over1} * {220\over b} =\] cross multiply